A rubber ball with a mass of 0.155 kg is dropped from rest. From what height (in m) was the ball dropped, if the magnitude of the ball's momentum is 0.710 kg · m/s just before it lands on the ground?
Answer #1

momentum = mass *velocity

p = mv

0.710 = 0.155*v

v = 4.58 m/s

KE = PE

0.5mv^2 = mgh

h = v^2 / 2g

= 4.58^2 / 2*9.8

= 1.07 m

so the height is 1.07 m